OpenStudy (chancemorris123):

Quadrilateral ABCD has side lengths of AB = 2z + 2, BC = 8, CD = 5z – 10, and AD = 8. Find the value of z for which quadrilateral ABCD is a parallelogram. A. 1 B. 2 C. 3 D. 4

OpenStudy (welshfella):

opposite sides of a parallelogram will be equal

OpenStudy (chancemorris123):

? can you help more plz :)

OpenStudy (welshfella):

so AB will be equal to CD

OpenStudy (chancemorris123):

oh so i + them?

OpenStudy (welshfella):

so 2z + 2 = 5z - 10 solve for z

OpenStudy (chancemorris123):

3z=12

OpenStudy (welshfella):

solve for z and you've got it

OpenStudy (chancemorris123):

then 3/12

OpenStudy (welshfella):

yes

OpenStudy (chancemorris123):

z=4

OpenStudy (welshfella):

right

OpenStudy (chancemorris123):

then what ?

OpenStudy (welshfella):

thats it the value you want is 4 check the options

OpenStudy (chancemorris123):

what? i dont get it

OpenStudy (welshfella):

they wanted to know the value of z which would make it a parallleogram opposite sides are equal thats why for it to be a paralleogram 2z +2 = 5z - 10 and this is true when z = 4

OpenStudy (welshfella):

so the answer is option D

OpenStudy (chancemorris123):

ohhhhhhhhhh ok thanks man thanks for the help

OpenStudy (welshfella):

ye

OpenStudy (chancemorris123):

chance morris out pyah